OpenStudy (anonymous):

FIND THE SLOPE: (1/2,2/3) AND (-2/3,3/4)

OpenStudy (lgbasallote):

(3/4 - 2/3)/(-2/3 - 1/2) (1/12)/(-7/6) simplifying... m = -1/14

OpenStudy (anonymous):

slope=y2-y1/x2-x1 =(3/4-2/3)/(-2/3-1/2) =1/12/-7/6 =-1/14

OpenStudy (anonymous):

but my teacher said that the answer is 1/7 i don't no why...?